# Towards Long-Horizon Vision-Language Navigation: Platform, Benchmark and Method

### v2

@@ 1 Introduction
Current Vision-Language Navigation (VLN) benchmarks and methods primarily focus on single-stage or short-term tasks, which involve simple objectives and limited action sequences, making them suitable for controlled settings but insufficient for real-world applications [21, 29, 22] (see Figure 1). In practical scenarios, agents must handle complex, long-horizon instructions that span multiple sub-tasks, requiring ongoing decision-making, dynamic re-planning, and sustained reasoning across extended periods [7, 40, 20, 27, 31]. These capabilities are crucial for applications like autonomous assistants or service robots where coherent navigation over a long temporal horizon is essential. To address this gap, we propose, for the first time, a new task-coded Long-Horizon VLN (LH-VLN), to evaluate and enhance agents’ abilities to manage multi-stage, context-rich navigation tasks that more accurately reflect real-world complexity.
@@
Methodology-wise, existing VLN methods typically rely on discretizing the environment into static points for path prediction, limiting adaptability in complex, dynamic settings [2, 49, 41, 24]. To bridge this gap and enhance real-world applicability in LH-VLN tasks, we introduce a Multi-Granularity Dynamic Memory (MGDM) module designed to enhance the model’s adaptability and memory handling. The MGDM module operates by integrating both short-term and long-term memory mechanisms. While short-term memory blurring and forgetting functions help the model focus on recent, relevant information, long-term memory retrieval pulls in key historical data from previous navigation steps [32]. This combination allows the model to adjust to environmental changes and retain context over extended sequences, addressing the challenges of sustained reasoning and adaptive re-planning in dynamic environments. By leveraging MGDM, our approach achieves new state-of-the-art performance on the LH-VLN task, demonstrating its effectiveness in maintaining coherent decision-making and robust navigation over long, multi-stage tasks.
Our contributions can be summarized as follows:
@@
We present the MGDM model, designed to enhance model adaptability in dynamic settings through combined short-term and long-term memory mechanisms, achieving outstanding performance on the LH-VLN task.
@@ 2.2 Benchmark for Vision-Language Navigation
The progression of VLN tasks has been propelled by a range of datasets, each introducing unique challenges and enhancing evaluation benchmarks for embodied agents performing tasks in complex environments. Early datasets, such as Room-to-Room (R2R) [3] and its extension Room-for-Room (R4R) [11], focus on step-by-step navigation through predefined paths with fine-grained instructions based on static images, while later datasets like VLN-CE [15] shift towards continuous navigation in dynamic spaces, requiring more flexible decision-making. More recent datasets, including CVDN [34], REVERIE [28], and SOON [50], further broaden the scope of VLN by integrating dialogue history, object localization, and complex instruction comprehension, pushing agents to understand high-level natural language commands and locate specific targets. Meanwhile, OVMM [44] and Behavior-1K [17] add layers of complexity by incorporating navigation, manipulation, and object interaction, simulating extended real-world tasks that involve multiple sub-tasks. Despite these progresses, there is still a notable gap in benchmarks that support LH-VLN with multi-stage sub-tasks in highly complex environments.
@@ 3.1.1 Forward Data Generation
In the forward data generation phase, we utilize GPT-4 to create task instructions by synthesizing scene assets and robot configurations, as shown in Fig. 2 (a). Specifically, our scene assets come from the HM3D dataset [42], which offers a rich collection of 3D panoramic scenes annotated semantically across 216 settings, providing an extensive foundation for task creation. Additionally, robot configurations are carefully tailored to different robotic platforms, such as Boston Dynamics’ Spot and Hello Robot’s Stretch, each with unique camera heights, task spaces, and sensor parameters to accommodate a variety of tasks.
@@
With these assets and configurations as the initial resource pool, a custom-designed prompt serves as the input for GPT-4, which combines scene details , robot configurations , and task goals . Then GPT-4 outputs an instruction list , including the sub-task and multi-stage instructions, and is denoted the GPT-4. This list is imported into the Habitat3 simulator , where an expert model or the well-trained navigation model guides the agent through the task, which the expert model is a navmesh model and greedy pathfinder algorithm built from Habitat [8, 16]. The simulator autonomously generates trajectories , the foundational data for subsequent splitting into task segments, as formulated:
@@ 3.2 The LHPR-VLN Benchmark
Our LHPR-VLN benchmark defines a complex task that includes multiple single-stage subtasks. For an LHPR-VLN task, the basic format is: “Find something somewhere, and take it to something somewhere, then…”. Each complex task involves locating an object at a specified initial location and transporting it to a designated target location, potentially encompassing two to four sequential navigation sub-tasks. The embodied agent needs to sequentially complete these single-stage navigation tasks to ultimately fulfill the instruction. For each single-stage navigation task, the agent must approach within a 1-meter geodesic distance of the target object, ensuring the object is positioned within a 60-degree horizontal field of view to maintain task fidelity.
Throughout navigation, the agent acquires observational data from three perspectives (, , ) and is permitted to execute fundamental actions: turn left, move forward, turn right, and stop. When the agent selects the “stop” action, the sub-task is deemed complete, and task success is evaluated based on the agent’s final positional state relative to the target. Table 1 presents a comparison between representative VLN benchmarks, our LHPR-VLN is the first LH-VLN benchmark, containing 3,260 multi-stage and step-by-step VLN tasks from 216 complex scenes, with an average of 150 task action steps and 18.17 instruction length.
@@ 3.3 Reasonable Metrics
where denotes the success of the -th sub-task.
@@ 4 Multi-Granularity Dynamic Memory Model
To achieve robust LH-VLN, our proposed Multi-Granularity Dynamic Memory (MGDM) model follows the general VLN pipeline and comprises three essential components: the base model, the Chain-of-Thought (CoT) Feedback module, and Adaptive Memory Integration and Update (AMIU), as shown in Fig. 4. These components enable robust performance in LH-VLN, addressing challenges related to spatial awareness, instruction comprehension, and task continuity across long-horizon sequences.
@@ 4.1 Base Model
The base framework is designed to align with the standard structure of VLN models. For scene observation, the model encodes multi-directional visual information using a pre-trained visual encoder vit. Each observed image is processed into visual features . To integrate scene information across multiple directions, a Transformer encoder is used for multi-view feature fusion. The directional image features are processed through the Transformer encoder, resulting in contextually enriched representations that capture inter-relational information across views.
@@ 4.2 Navigate with CoT and Memory
To address the limited interpretability and susceptibility to “hallucinations” [35] in LLM-based VLN models (wherein the agent completes tasks without true comprehension), we introduce a Chain-of-Thought (CoT) [39] Feedback module that receives task instructions and, based on historical observation of corresponding memory, generates a chain of thought and constructs language prompts. This module aims to enhance the agent’s reasoning capability by iteratively refining its task understanding and action planning.
@@
When the memory length reaches a set maximum , dynamic forgetting is triggered. Each memory element has an associated confidence score , representing the model’s confidence in action . The memory sequence thus has an associated confidence vector .
The forgetting module employs a “pooling function” that we define it as . represents the pooling operation with a window size of 2 applied to the element and its neighboring elements in :
@@
Long-term memory serves as a reinforcement mechanism. As the agent navigates, long-term memory retrieves relevant observations and actions from the dataset, matching them with the agent’s current context to provide guidance. The retrieval process selects the top matching observation-action pairs, which are weighted to inform the current decision vector. This memory is sourced from the LHPR-VLN dataset, reinforcing prior learning:
Thus, the final instance can be formulated as:
The action decision is computed by averaging the retrieved actions:
@@ 5.1 Experimental Settings
Sensors: For each action step, the agent receives RGB observations from there directions of front, left (+60°), and right (-60°). Depth images for these three directions can also be customized
@@
Training Settings: We alternately use imitation learning and trajectory-based supervised learning for training. The LLM is Vicuna 7B v0 [4], and the visual encoder is the ViT model from EVA-CLIP-02-Large [33]. The visual encoder remains frozen during training. Through out the training phase, we utilize the Adam optimizer with a learning rate of 3e-5.
Metrics: In addition to the LH-VLN metrics ISR, CSR , and CGT, we conducted tests based on benchmark metrics for navigation tasks [3], including SR (Success Rate), SPL (Success weighted by Path Length), OSR (Oracle Success Rate), and NE (Navigation Error). For SR, OSR and SPL, the task is considered successful only when all sub-tasks in a LH-VLN task are completed correctly in the logical sequence of instructions. For NE, only when the agent takes the action of ‘stop’, the NE counts.
5.1.1 Baseline Models
@@ 5.2 Result Analysis
RQ3: Furthermore, all models perform better in terms of ISR, CSR, and CGT on LH-VLN tasks with 3-4 subtasks than on those with 2-3 subtasks. This may be due to the fact that while longer and multi-stage tasks may be more difficult, the memory accumulated from previous stages can help the VLN model complete subtasks in subsequent stages. This may suggest the significance of developing VLN models for multi-stage complex tasks.
It is worth noting that our MGDM has a relatively low NE. when tasks are so difficult that the model performs poorly, NE reflects the gap between the model performance and success. This suggests that our MGDM may have greater potential for LH-VLN. Additionally, in the step-by-step tasks shown in Table 3, although our MGDM has higher OSR and lower NE, its SR and SPL metrics are both 0. This indicates that our MGDM currently faces an issue in effectively determining whether the goal has been achieved.
@@ 5.3 Ablation Studies
We performed ablation studies on multi-granularity dynamic memory module and the chain-of-thought (CoT) feedback module, with results shown in Table 4. As observed, the model’s performance is significantly affected whether the CoT feedback module or the multi-granularity dynamic memory module is ablated. This indicates the crucial role of chain-of-thought generation and memory in the model’s ability to solve LH-VLN tasks. From the perspective of NE, especially the multi-granularity dynamic memory module, it has significant impact on model’s performance. This is also reflected in the visualization analysis of a successful long-horizon navigation example (see Figure 5). The agent’s actions are very chaotic at the beginning (1-3 steps). It only acts effectively once the memory sequence reaches a certain length. This further underscores the importance of memory module design for LH-VLN tasks.
@@ 8.3 Cases Study
For Case A, the agent quickly found the living room and identified the item on the table in Observation 3 as the target. However, in Observation 4, it determined that it was not a ”bag”, prompting a strategy change and further exploration of the scene (Observations 5–12). After completing the scene exploration, the agent resumed the search for the target.
For Case B, the agent was initially unable to determine the current scene and decided to rotate in place (Observations 1–4). It then tentatively searched for the living room. Upon realizing the current direction did not lead to the living room, it turned around (Observations 6–7) and moved toward the living room area (Observations 8–10). Since it did not find the ”device” in the living room, the agent chose to explore another direction (Observations 11–12).
@@ 9.4 More experimental results
Additionally, in tasks with the Spot robot configuration, MGDM’s performance is slightly lower than NaviLLM+GPT-4. Notably, for tests involving NaviLLM-related models, their performance shows a significant advantage under the Spot robot configuration. This suggests that the NaviLLM model is relatively better suited for Spot-related tasks. This advantage may stem from NaviLLM’s pretraining data being more closely aligned with the Spot robot configuration, giving it a certain edge in this aspect. In contrast, MGDM, due to the design of its memory module, is more likely to ”remember” small objects in the scene. This may reduce the performance gap between the Spot and Stretch robot configurations, leading to more balanced results across both setups.

### v3

@@ 1 Introduction
Current Vision-Language Navigation (VLN) benchmarks and methods primarily focus on single-stage or short-term tasks, which involve simple objectives and limited action sequences, making them suitable for controlled settings but insufficient for real-world applications [24] (see Figure LABEL:fig:intro). In practical scenarios, agents must handle complex, long-horizon instructions that span multiple sub-tasks, requiring ongoing decision-making, dynamic re-planning, and sustained reasoning across extended periods [8, 30, 33]. These capabilities are crucial for applications like autonomous assistants or service robots where coherent navigation over a long temporal horizon is essential. To address this gap, we propose, for the first time, a new task-coded Long-Horizon VLN (LH-VLN), to evaluate and enhance agents’ abilities to manage multi-stage, context-rich navigation tasks that more accurately reflect real-world complexity.
@@
Existing VLN methods typically rely on discretizing the environment into static points for path prediction, limiting adaptability in complex, dynamic settings [2, 50, 42, 26]. To bridge this gap and enhance real-world applicability in LH-VLN tasks, we introduce a Multi-Granularity Dynamic Memory (MGDM) module to enhance the model’s adaptability and memory handling. The MGDM module operates by integrating both short-term and long-term memory mechanisms. While short-term memory blurring and forgetting functions help the model focus on recent, relevant information, long-term memory retrieval pulls in key historical data from previous navigation steps [34]. This combination allows the model to adjust to environmental changes and retain context over extended sequences, addressing the challenges of sustained reasoning and adaptive re-planning in dynamic environments. With MGDM, we achieve state-of-the-art performance on the LH-VLN task, demonstrating its effectiveness in maintaining coherent decision-making and robust navigation over long, multi-stage tasks. Our contributions can be summarized as follows:
@@
We present the MGDM model, designed to enhance model adaptability in dynamic settings through combined short-term and long-term memory mechanisms.
@@ 2.2 Benchmark for Vision-Language Navigation
The progression of VLN tasks has been propelled by a range of datasets, each introducing unique challenges and enhancing evaluation benchmarks for embodied agents performing tasks in complex environments. Early datasets, such as Room-to-Room (R2R) [3] and its extension Room-for-Room (R4R) [12], focus on step-by-step navigation through predefined paths with fine-grained instructions based on static images, while later datasets like VLN-CE [18] shift towards continuous navigation in dynamic spaces, requiring more flexible decision-making. More recent datasets, including CVDN [36], REVERIE [31], and SOON [51], further broaden the scope of VLN by integrating dialogue history, object localization, and complex instruction comprehension, pushing agents to understand high-level natural language commands and locate specific targets. Meanwhile, OVMM [45] and Behavior-1K [21] add layers of complexity by incorporating navigation, manipulation, and object interaction, simulating extended real-world tasks that involve multiple sub-tasks. IVLN [19] and Goat-Bench [15] allow the agent to continuously complete multiple independent single-target navigation tasks while maintaining memory. Despite these progresses, there is still a notable gap in benchmarks that support LH-VLN with multi-stage sub-tasks in highly complex environments.
@@ 3.1.1 Forward Data Generation
In the forward data generation phase, we utilize GPT-4 to create task instructions by synthesizing scene assets and robot configurations, as shown in Fig. 1. Specifically, our scene assets come from the HM3D dataset [43], which offers a rich collection of 3D panoramic scenes annotated semantically across 216 settings, providing an extensive foundation for task creation. Additionally, robot configurations are carefully tailored to different robotic platforms, such as Boston Dynamics’ Spot and Hello Robot’s Stretch, each with unique camera heights, task spaces, and sensor parameters to accommodate a variety of tasks.
@@
| IVLN [19] | - | - | M3D&Habitat | Iterative VLN | 72 | 789 |
| Goat-Bench [15] | - | - | Habitat | Iterative VLN | 181 | 725360 |
@@
With these assets and configurations as the initial resource pool, a custom-designed prompt serves as the input for GPT-4, which combines scene details and robot configurations . Then GPT-4 outputs an instruction list , including the sub-task and multi-stage instructions, and is denoted the GPT-4. This list is imported into the Habitat3 simulator , where an expert model or a well-trained navigation model guides the agent through the task, which the expert model is a navmesh model and greedy pathfinder algorithm built from Habitat [9, 20]. The simulator autonomously generates trajectories , the foundational data for subsequent splitting into task segments:
@@ 3.2 The LHPR-VLN Benchmark
Our LHPR-VLN benchmark defines a complex task that includes multiple single-stage subtasks. For an LHPR-VLN task, the basic format is: “Find something somewhere, and take it to something somewhere, then…". Each complex task involves locating an object at a specified initial location and transporting it to a designated target location, potentially encompassing two to four sequential navigation sub-tasks. The embodied agent needs to sequentially complete these single-stage navigation tasks to ultimately fulfill the instruction. For each single-stage navigation task, the agent must approach within a 1-meter geodesic distance of the target object, ensuring the object is positioned within a 60-degree horizontal field of view to maintain task fidelity.
Throughout navigation, the agent acquires observational data from three perspectives (, , ) and is permitted to execute fundamental actions: turn left, move forward, turn right, and stop. When the agent selects the “stop" action, the sub-task is deemed complete, and task success is evaluated based on the agent’s final positional state relative to the target. Table 6 presents a comparison between representative VLN benchmarks, our LHPR-VLN is the first LH-VLN benchmark, containing 3,260 multi-stage and step-by-step VLN tasks from 216 complex scenes, with an average of 150 task action steps and 18.17 instruction length.
@@ 3.3 Reasonable Metrics
where denotes the success of the -th sub-task in .
@@
We also designed a metric Target Approach Rate (TAR) based on NE to reflect the model’s performance in cases where the navigation success rate is relatively low. The relevant settings can be found in the supplementary materials.
@@ 4 Multi-Granularity Dynamic Memory Model
To achieve robust LH-VLN, our Multi-Granularity Dynamic Memory (MGDM) model follows the general VLN pipeline and comprises three essential components: the base model, the Chain-of-Thought (CoT) Feedback module, and Adaptive Memory Integration and Update (AMIU), as shown in Fig. 3. These components enable robust performance in LH-VLN, addressing challenges related to spatial awareness [27], instruction comprehension [4], and task continuity [13] across long-horizon sequences.
@@ 4.1 Base Model
The base model aligns with the standard structure of VLN models. For scene observation, the model encodes multi-directional visual information using a pre-trained visual encoder vit. Each observed image is processed into visual features . To integrate scene information across multiple directions, a Transformer encoder is used for multi-view feature fusion. The directional image features are processed through the Transformer encoder, resulting in contextually enriched representations that capture inter-relational information across views.
@@ 4.2 Navigate with CoT and Memory
To address the limited interpretability and susceptibility to “hallucinations" [37] in LLM-based VLN models (wherein the agent completes tasks without true comprehension), we introduce a Chain-of-Thought (CoT) [41] Feedback module that receives task instructions and, based on historical observation of corresponding memory, generates a chain of thought and constructs language prompts. This module aims to enhance the agent’s reasoning capability by iteratively refining its task understanding and action planning.
@@
When the memory length reaches a set maximum , dynamic forgetting is triggered. Each memory element has an associated confidence score , representing the model’s confidence in corresponding action. The memory sequence thus has an associated confidence vector .
The forgetting module employs a “pooling function" that we define it as . represents the pooling operation with a window size of 2 applied to the element and its neighboring elements in , which reduces its length by one:
@@
Long-term memory serves as a reinforcement mechanism. As the agent navigates, long-term memory retrieves relevant observations and actions based on target from the dataset, matching them with the agent’s current observation to provide guidance. The retrieval process selects the top matching observation-action pairs, which are weighted to inform the current decision vector. This memory is sourced from the LHPR-VLN dataset, reinforcing prior learning:
Thus, the indices of the selected can be formulated as:
The action decision is weighted by averaging the retrieved actions:
@@ 5.1 Experimental Settings
Sensors: For each action step, the agent receives RGB observations from there directions of front, left (+60°), and right (-60°). Depth images for these three directions can also be customized.
@@
Training Settings: We alternately use imitation learning and trajectory-based supervised learning. The LLM is Vicuna 7B v0 [5], and the visual encoder is the ViT model from EVA-CLIP-02-Large [35]. The visual encoder remains frozen during training. In the training phase, we utilize the Adam optimizer with a learning rate of 3e-5.
Metrics: Besides our metrics ISR, CSR , and CGT, we also used traditional metrics [3], including SR (Success Rate), SPL (Success weighted by Path Length), OSR (Oracle Success Rate), and NE (Navigation Error). For SR, OSR and SPL, the task is considered successful only when all sub-tasks in a LH-VLN task are completed correctly in the logical sequence of instructions. For NE, only when the agent takes the action of ‘stop’, the NE counts.
5.2 Baseline Models
@@ 5.3 Result Analysis
RQ3: Furthermore, all models perform better in ISR, CSR, and CGT on LH-VLN tasks with 3-4 subtasks than on those with 2-3 subtasks. This may be due to the fact that while longer and multi-stage tasks may be more difficult, the memory accumulated from previous stages can help the VLN model complete subtasks in subsequent stages. This may suggest the significance of developing VLN models for multi-stage complex tasks. The tendency of navigation target distribution in the LH-VLN task with different numbers of subtasks and task settings may also influence this result. Relevant details can be found in the supplementary materials. It is worth noting that our MGDM has a relatively low NE. when tasks are so difficult that the model performs poorly, NE reflects the gap between the model performance and success. This suggests that our MGDM may have greater potential for LH-VLN. Additionally, in the step-by-step tasks shown in Table 3, although our MGDM has higher OSR and lower NE, its SR and SPL metrics are both 0. This indicates that our MGDM faces an issue in effectively determining whether the goal has been achieved.
@@ 5.4 Ablation Studies
We performed ablation studies on multi-granularity dynamic memory module, long term memory module and the chain-of-thought (CoT) feedback module, with results shown in Table 4. As observed, the model’s performance is significantly affected whether the CoT feedback module, long term memory module or the multi-granularity dynamic memory module is ablated. This indicates the crucial role of chain-of-thought generation and memory in the model’s ability to solve LH-VLN tasks. From the perspective of NE, especially the multi-granularity dynamic memory module, it has significant impact on model’s performance. This is also reflected in the visualization analysis of a successful long-horizon navigation example (see Figure 4). The agent’s actions are very chaotic at the beginning (1-3 steps). It only acts effectively once the memory sequence reaches a certain length. This further underscores the importance of memory module design for LH-VLN tasks.
@@ 7 Symbol Table
| The number of viewpoints in the observation. | |
| The indices of the selected elements | |
@@ 8.2 Dataset Statistics
We also conducted statistical analysis on more detailed data regarding task goals at different stages in the dataset, as shown in Table. 6.
| Type | Task Steps | Nav Dis | Obj Num | Floor Span | Reentry Rate | Area Association |
| 2 subtasks | 68.39 | 11.60, 11.23, -, - | 288.43 | 1.05 | 27.44% | 1.02 |
| 3 subtasks | 53.30 | 10.32, 11.47, 4.74, - | 357.05 | 1.44 | 42.11% | 1.89 |
| 4 subtasks | 51.23 | 9.75, 10.45, 4.14, 9.39 | 273.37 | 1.88 | 48.84% | 2.12 |
8.3 Extra Metric
We designed a metric Target Approach Rate (TAR) based on NE to reflect the model’s performance in cases where the navigation success rate is relatively low. For the -th subtask of the -th task, we calculate the Target Approach Rate (TAR):
where is NE of the -th subtask of the -th task, is the distance to be considered success, is ground truth of the -th subtask of the -th task.
@@ 8.4 Cases Study
For Case A, the agent quickly found the living room and identified the item on the table in Observation 3 as the target. However, in Observation 4, it determined that it was not a "bag", prompting a strategy change and further exploration of the scene (Observations 5–12). After completing the scene exploration, the agent resumed the search for the target.
For Case B, the agent was initially unable to determine the current scene and decided to rotate in place (Observations 1–4). It then tentatively searched for the living room. Upon realizing the current direction did not lead to the living room, it turned around (Observations 6–7) and moved toward the living room area (Observations 8–10). Since it did not find the "device" in the living room, the agent chose to explore another direction (Observations 11–12).
@@ 9.4 More experimental results
Additionally, in tasks with the Spot robot configuration, MGDM’s performance is slightly lower than NaviLLM+GPT-4. Notably, for tests involving NaviLLM-related models, their performance shows a significant advantage under the Spot robot configuration. This suggests that the NaviLLM model is relatively better suited for Spot-related tasks. This advantage may stem from NaviLLM’s pretraining data being more closely aligned with the Spot robot configuration, giving it a certain edge in this aspect. In contrast, MGDM, due to the design of its memory module, is more likely to "remember" small objects in the scene. This may reduce the performance gap between the Spot and Stretch robot configurations, leading to more balanced results across both setups.
@@
For the issue where tasks with fewer subtasks yield worse results, we consider the following reasons:
The first reason that the model performs worse in shorter tasks is that the model needs sufficient steps to warm up, i.e., setting up correct initial direction and locate targets. Though shorter tasks may seem easier, they leave fewer steps after initialization, making it more challenging for the agent to complete the task, thus leading to worse performance.
In our dataset analysis, as the number of subtasks in complex tasks increases, the average ground truth steps per subtask decrease (68.39, 53.30, 51.23), while the average maximum number of identical regions in the task increases (1.02, 1.89, 2.12), the probability of the agent observing critical information about subsequent subtasks during current subtasks increases(27.44%, 42.11%, 48.84%). This suggests that with more subtasks, the probability of different subtask targets appearing in the same region increases, the average number of steps required for each subtask decreases, and the average difficulty of each subtask decreases.
This is also reflected in the distances between subtask goals. In Table. 6, the navigation distance for the third subtask (i.e., the distance between the second and third goals) is always significantly smaller than the navigation distances at other stages. (4.74, 4.14 to 10.32, 11.47 and 9.74, 10.45) This indicates that in complex tasks with three or more subtasks, once the second goal is found, it becomes much easier to locate the third goal.